# Parallelizing MCMC Across the Sequence Length

### v1

@@ Abstract
These authors contributed equally.
1 Department of Statistics, Stanford University, Stanford, CA, USA
2 Wu Tsai Neurosciences Institute, Stanford University, Stanford, CA, USA
3 Thomas J. Watson Research Center, IBM Research, Yorktown Heights, NY, USA
Correspondence should be addressed to D.M.Z. (dzoltowski@stanford.edu), S.W. (skylerw@stanford.edu), and S.W.L. (scott.linderman@stanford.edu).
@@ 1 Introduction
Markov chain Monte Carlo (MCMC) algorithms generate samples from a target distribution via a sequential process, incurring a computational cost that is linear in the number of samples (Diaconis, 2009; Geyer, 2011; Gelman et al., 2013). While modern hardware like GPUs and TPUs can yield dramatic improvements in processing speed for algorithms that can be parallelized, capitalizing on these hardware advances remains a challenge for MCMC. In particular, while parallel resources can be used to simulate many independent chains in parallel (Sountsov et al., 2024; Margossian et al., 2024), the runtime remains linear in the sequence length.
@@
In this paper, we present a promising approach for parallelizing MCMC algorithms across the chain length to obtain sublinear chain-length complexity. The core idea is to adapt methods for the parallel evaluation of nonlinear recursions (Danieli et al., 2023; Lim et al., 2024; Gonzalez et al., 2024) to MCMC algorithms. These methods formulate the state sequence of a nonlinear sequence model as the solution of an optimization problem and iteratively solve for the state sequence via a parallel-in-time formulation of Newton’s method.
@@
We show that these methods can parallelize widely used classes of MCMC samplers across the chain length, including Gibbs sampling (Geman and Geman, 1984), the Metropolis-adjusted Langevin algorithm (MALA) (Besag, 1994), and Hamiltonian Monte Carlo (HMC) (Duane et al., 1987; Neal et al., 2011; Betancourt, 2017). Across these examples, we often find that dozens of parallel evaluations are sufficient to generate anywhere from thousands to hundreds of thousands of samples. We also propose two new variations of the parallel Newton method that reduce computational costs and improve efficiency for certain MCMC applications. With these approaches, we find that the proposed parallel algorithms can dramatically accelerate MCMC sampling.
@@
The paper is organized as follows. First, we provide background on using Newton’s method to parallelize the evaluation of nonlinear recursions. Then we describe how this approach can be used to parallelize Gibbs sampling, MALA, and HMC across the sequence length. For HMC, we additionally propose to parallelize only the leapfrog integration within each proposal. We then describe innovations that improve the scalability of parallel Newton methods and devise a novel block quasi-Newton method adapted for parallelizing leapfrog integration. Finally, we present experiments across multiple MCMC algorithms that characterize the number of parallel iterations to convergence and wall-clock times of the proposed algorithms. In Figure˜1 we illustrate this approach for simulating 100K samples of HMC targeting a 2D distribution with high-curvature. The parallel HMC algorithm converges to the ground-truth sequential trace in only 147 steps. Our implementation is available at https://github.com/lindermanlab/parallel-mcmc.
@@
Figure 1: Parallel evaluation of 100K HMC samples targeting the Rosenbrock distribution. (A) Dark lines (black for sequential, blue for parallel) show simulated samples of the second coordinate for the first 1000 samples at three iterations of either sequential or parallel MCMC. The light gray line is the final sequential chain. (B) 100K samples after 1, 10, 25, or 147 parallel iterations (blue), with convergence to the sequential trace (gray) at iteration 147.
@@ 2 Background
[…]
In contrast, nonlinear sequential processes such as recurrent neural networks and, for our purposes, common MCMC algorithms, are not directly parallelizable and instead are typically evaluated sequentially. However, Danieli et al. (2023) and Lim et al. (2024) showed that we can view the states of a nonlinear recursion as the solution of a fixed-point equation, and that a parallelized form of Newton’s method can be used to solve for the state sequence. Lim et al. (2024) and Danieli et al. (2023) observed that iteratively linearizing the transition functions, $f_{t}$, around a guess for the state trajectory, $\mathbf{s}_{1:T}^{(i)}$, and evaluating the resulting linear system to obtain a new trajectory, $\mathbf{s}_{1:T}^{(i+1)}$, is equivalent to using Newton’s method on an appropriately defined residual. More precisely, given $\mathbf{s}_{0}$ and $f_{1:T}$, the residual is the vector of temporal differences
@@
with dynamics given by the time-varying Jacobians, $\boldsymbol{J}_{t}$, and inputs, $\mathbf{u}_{t}$. The Jacobians and inputs solely depend on the states from the previous Newton iteration, and therefore they can be computed in parallel. Crucially, this linear dynamical system can also be solved in $\mathcal{O}(\log T)$ time on a parallel machine using the parallel scan algorithm described above. This parallel update is repeated until the state sequence converges within a specified numerical tolerance $\delta$. Lim et al. (2024) called this the DEER algorithm, and Danieli et al. (2023) called it DeepPCR. To avoid confusion, we will refer to this approach as DEER or simply Newton’s method.
@@
The DEER algorithm requires computing, storing, and multiplying $T$ Jacobian matrices $\boldsymbol{J}_{t}$ of size $D\times D$. Therefore, DEER must perform $\mathcal{O}(TD^{3})$ work and requires $\mathcal{O}(TD^{2})$ memory, which can be prohibitive for scaling to larger dimensionalities. To address this issue, Gonzalez et al. (2024) proposed using quasi-Newton methods to reduce computational complexity. The proposed quasi-DEER method retains only the diagonal of the Jacobian matrices, $\mathrm{diag}(\boldsymbol{J}_{t})$, and evaluates the following linear recursion
$\mathbf{s}_{t}^{(i+1)}=\mathrm{diag}(\boldsymbol{J}_{t})\,\mathbf{s}_{t-1}^{(i+1)}+f_{t}(\mathbf{s}_{t-1}^{(i)})-\mathrm{diag}(\boldsymbol{J}_{t})\,\mathbf{s}_{t-1}^{(i)}.$ (4)

### v2

@@ 1 Introduction
Markov chain Monte Carlo (MCMC) algorithms sequentially generate samples from a target distribution, with computational cost linear in the length of the chain [1, 2, 3]. While modern hardware like GPUs and TPUs can dramatically accelerate parallelizable algorithms, leveraging these advances for MCMC remains challenging. In particular, while parallel resources can be used to simulate many independent chains in parallel [4, 5], the runtime remains linear in the sequence length.
@@
In this paper, we present a promising approach for parallelizing MCMC algorithms across the sequence length to obtain sublinear chain-length complexity. The core idea is to adapt methods for the parallel evaluation of nonlinear recursions [6, 7, 8] to MCMC algorithms. These methods formulate the state sequence of a nonlinear sequence model as the solution of an optimization problem and iteratively solve for the state sequence via a parallel-in-time formulation of Newton’s method.
@@
We show that these methods can parallelize widely used classes of MCMC samplers across the chain length, including Gibbs sampling [9], the Metropolis-adjusted Langevin algorithm (MALA) [10], and Hamiltonian Monte Carlo (HMC) [11, 12, 13]. Across these examples, we often find that dozens of parallel evaluations are sufficient to generate anywhere from thousands to hundreds of thousands of samples. We also propose two new variations of the parallel Newton method that reduce computational costs and improve efficiency for certain MCMC applications. With these approaches, we find that the proposed parallel algorithms can dramatically accelerate MCMC sampling.
@@
The paper is organized as follows. We first review how Newton’s method can parallelize nonlinear recursions, then show how this framework applies to Gibbs sampling, MALA, and HMC across the sequence length. For HMC, we additionally propose parallelizing only the leapfrog integration within each proposal. Next, we introduce scalability improvements to parallel Newton methods and a novel block quasi-Newton variant adapted for parallelizing leapfrog integration. Finally, we present experiments demonstrating wall-clock speedups of parallel MCMC sampling across multiple examples. In Figure˜1, we illustrate our approach by simulating 100K HMC samples from a high-curvature 2D distribution, where parallel HMC converges to the sequential trace in just 147 steps. Our implementation is available at https://github.com/lindermanlab/parallel-mcmc.
@@
Figure 1: Parallel evaluation of 100K HMC samples targeting the Rosenbrock distribution. (A) Dark lines (black: sequential, blue: parallel) show simulated samples of the second coordinate for the first 1K samples at three iterations of each method. The light gray line is the final sequential chain. (B) 100K samples after 1, 10, 25, or 147 parallel iterations (blue), converging to the sequential trace (gray) by iteration 147.
@@ 2 Background
Our work builds on and extends previous approaches for parallelizing linear and nonlinear recursions. Consider a recursion, $\mathbf{s}_{t}=f_{t}(\mathbf{s}_{t-1})$, that when given an initial state $\mathbf{s}_{0}$ yields a sequence of states, $\mathbf{s}_{1:T}=(\mathbf{s}_{1},\ldots,\mathbf{s}_{T})\subset\mathbb{R}^{D}$, according to the transition functions, $f_{t}:\mathbb{R}^{D}\mapsto\mathbb{R}^{D}$. The obvious way to solve for the sequence of states is to evaluate the recursion sequentially. However, if $f_{t}$ is an affine function (i.e., $f_{t}(\mathbf{s}_{t-1})=\boldsymbol{J}_{t}\mathbf{s}_{t}+\mathbf{u}_{t}$ for some matrix $\boldsymbol{J}_{t}$ and vector $\mathbf{u}_{t}$), then the system can be evaluated in $\mathcal{O}(\log T)$ time on a parallel machine using a parallel (a.k.a. associative) scan [14]. The key insight is that a composition of updates, $f_{t+1}\circ f_{t}$, is still an affine function. With $\mathcal{O}(T)$ processors, one can compute all pairs of updates in parallel, yielding a sequence that is half as long. By repeating this process $\mathcal{O}(\log T)$ times, one can obtain the entire sequence in sublinear time. This parallel algorithm underlies many modern machine learning models for sequential data [15, 16, 17].
In contrast, nonlinear sequential processes such as recurrent neural networks and, for our purposes, common MCMC algorithms, are not directly parallelizable and instead are typically evaluated sequentially. However, Danieli et al. [6] and Lim et al. [7] showed that we can view the states of a nonlinear recursion as the solution of a fixed-point equation, and that a parallelized form of Newton’s method can be used to solve for the state sequence. Lim et al. [7] and Danieli et al. [6] observed that iteratively linearizing the transition functions, $f_{t}$, around a guess for the state trajectory, $\mathbf{s}_{1:T}^{(i)}$, and evaluating the resulting linear system to obtain a new trajectory, $\mathbf{s}_{1:T}^{(i+1)}$, is equivalent to using Newton’s method on an appropriately defined residual. More precisely, given $\mathbf{s}_{0}$ and $f_{1:T}$, the residual is the vector of temporal differences
@@
with dynamics given by the time-varying Jacobians, $\boldsymbol{J}_{t}$, and inputs, $\mathbf{u}_{t}$. The Jacobians and inputs only depend on the states from the previous Newton iteration, and thus can be computed in parallel. Crucially, this linear dynamical system can also be solved in $\mathcal{O}(\log T)$ time on a parallel machine using the parallel scan algorithm described above. This parallel update is repeated until the state sequence converges within a numerical tolerance $\delta$. This algorithm was termed “DEER” by Lim et al. [7] and “DeepPCR” by Danieli et al. [6]. For clarity, we refer to it as “DEER” or “Newton’s method.”
@@
The DEER algorithm requires computing, storing, and multiplying $T$ Jacobian matrices $\boldsymbol{J}_{t}$ of size $D\times D$. Therefore, DEER must perform $\mathcal{O}(TD^{3})$ work and requires $\mathcal{O}(TD^{2})$ memory, which can be prohibitive in larger dimensionalities. To reduce computational complexity, Gonzalez et al. [8] proposed quasi-DEER, a quasi-Newton method that retains only the diagonal of the Jacobian matrices, $\mathrm{diag}(\boldsymbol{J}_{t})$, and evaluates the following linear recursion
$\mathbf{s}_{t}^{(i+1)}=\mathrm{diag}(\boldsymbol{J}_{t})\,\mathbf{s}_{t-1}^{(i+1)}+f_{t}(\mathbf{s}_{t-1}^{(i)})-\mathrm{diag}(\boldsymbol{J}_{t})\,\mathbf{s}_{t-1}^{(i)}.\vskip-2.5pt$ (4)